### v1

# Virasoro Symmetry in Neural Network Field Theories

Abstract
Neural Network Field Theories (NN-FTs) can realize global conformal symmetries via embedding space architectures. These models describe Generalized Free Fields (GFFs) in the infinite width limit. However, they typically lack a local stress-energy tensor satisfying conformal Ward identities. This presents an obstruction to realizing infinite-dimensional, local conformal symmetry typifying 2d Conformal Field Theories (CFTs). We present the first construction of an NN-FT that encodes the full Virasoro symmetry of a 2d CFT. We formulate a neural free boson theory with a local stress tensor $T(z)$ by properly choosing the architecture and prior distribution of network parameters. We verify the analytical results through numerical simulation; computing the central charge and the scaling dimensions of vertex operators. We then construct an NN realization of a Majorana Fermion and an $\mathcal{N}=(1,1)$ scalar multiplet, which then enables an extension of the formalism to include super-Virasoro symmetry. Finally, we extend the framework by constructing boundary NN-FTs that preserve (super-)conformal symmetry via the method of images.
Introduction.— The interface between theoretical physics and machine learning Carleo:2019ptp has revealed a deep correspondence between Neural Networks (NNs) and Quantum Field Theory (QFT). In the infinite width limit, the output of a randomly initialized network converges to a Gaussian Process governed by a kernel $K(x,y)$ Neal:1996 ; NIPS1996_ae5e3ce4 ; Lee:2017wiy , effectively defining a free field theory Halverson:2020trp ; Halverson:2021aot . Finite-width corrections induce non-Gaussian interactions, providing a framework to simulate renormalization group flows and neural effective actions Dyer:2019uzd ; Yaida:2020 ; Roberts:2021principles ; Erdmenger:2022tqi ; Grosvenor:2022edge ; Schoenholz:2016 .
Central to the development of this program is building NNs that encode spacetime symmetries. While global conformal invariance ($SO(d+1,1)$) has been achieved via embedding space formalism Halverson:2024axc and extended to include conformal defects Billo:2016cpy ; Capuozzo:2025ozt , these constructions typically yield GFFs. While respecting global conformal symmetries, for generic scaling dimensions, these GFFs lack a local, conserved stress-energy tensor $T_{\mu\nu}$ satisfying conformal Ward identities. For NN-FTs with arbitrary dimensional input space, this would present an obstruction to the standard computation of Weyl anomalies but would still allow for the computation of correlation functions and OPEs. In two dimensions (2d), though, this is a fatal deficit: without a local $T(z)$, we cannot construct Virasoro generators $L_{n}$, and thus no local conformal symmetry. The lack of local conformal symmetry in 2d presents a significant barrier to the application of the NN-FT formalism to the study of, e.g., low dimensional critical phenomena and string theory.
In this Letter, we construct a neural architecture—the Log-Kernel Network (LK)—that supports a local stress tensor and demonstrate the emergence of the Virasoro algebra in NN-FTs. We validate this construction by computing the central charge $c=1$ and the spectrum of vertex operators for a theory of 2d free bosons. We then construct the Neural Majorana Fermion (NMF) and demonstrate the emergence of the super-Virasoro algebra. Finally, we demonstrate the robustness of the framework by identifying conformal boundary conditions in bosonic and fermionic NN-FTs.
The LK Architecture.— To realize a local stress tensor $T(z)\sim:(\partial\phi)^{2}:$, the underlying architecture $\phi$ that we are engineering as a 2d free boson must possess a log two-point function $\langle\phi(z)\phi(0)\rangle\sim-\ln|z|^{2}$ DiFrancesco:1997nk . We construct $\phi(x)$ on $\mathbb{R}^{2}\cong\mathbb{C}$ as a superposition of random Fourier features:
$\phi(x)=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}A_{j}\cos(\vec{k}_{j}\cdot\vec{x}+\gamma_{j})\,$, (1)
where phases $\gamma_{j}$ are drawn uniformly from $0,2\pi)$ and amplitudes $A_{j}=1$ are fixed. Here we note that any single realization of the network $\phi(x;\Theta)$ does not respect translation invariance, the ensemble does. The uniform integration over phases $\gamma_{j}$ enforces functional dependence only on separations, e.g. $\langle\phi(x)\phi(y)\rangle=f(x-y)$. Furthermore, although the Log-correlated field $\phi(x)$ shifts under special conformal transformations, its gradients $\partial\phi$ and Neural Vertex Operators (NVOs) $\mathcal{V}_{\alpha}$ transform covariantly. The crucial ingredient to restoring global conformal invariance is the distribution of wavevectors $\vec{k}_{j}$. To ensure scale invariance of the Gaussian Process kernel, the probability measure $d\mu(k)=p(k)d^{2}k$ must be invariant under dilatations $k\to\lambda k$. This fixes the spectral density to be $p(k)\propto|k|^{-2}$ in the infinite width limit.
To define a normalizable probability distribution, we restrict the support to an annulus $\Lambda_{IR}\leq|k|\leq\Lambda_{UV}$. The normalized density is:
$p(k)=\frac{1}{\mathcal{Z}}\frac{1}{|k|^{2}}\mathbb{I}_{\mathcal{D}}(k)\,,\quad\mathcal{Z}=2\pi\ln\left(\frac{\Lambda_{UV}}{\Lambda_{IR}}\right)\,.$ (2)
In the infinite-width limit, the kernel is the Fourier transform of this spectral density. Evaluating the integral yields the required logarithmic propagator:
$K(z,w)$ $\equiv\mathbb{E}[\phi(z)\phi(w)]=\frac{1}{2}\int d^{2}k\,p(k)e^{ik\cdot(z-w)}$
$\approx-\frac{1}{2\mathcal{Z}}(2\pi\ln|z-w|+\dots)=-C_{\phi\phi}\ln|z-w|\,.$ (3)
where $C_{\phi\phi}=\pi/\mathcal{Z}$ is a normalization constant determined by the variance of the weights.
We pause here to note the distinction between explicit symmetry breaking and regularization in the LK architecture. The spectral cutoffs $\Lambda_{UV}$ and $\Lambda_{IR}$ serve strictly as regulators for the probability measure. In the physical window $\Lambda_{UV}^{-1}\ll|z-w|\ll\Lambda_{IR}^{-1}$, the kernel scales as $K(\lambda z,\lambda w)\sim K(z,w)-\ln|\lambda|^{2}$, which corroborates the identification as an NN version of a 2d free boson CFT. The cutoff dependence is absorbed entirely into the field normalization constant $C_{\phi\phi}$. In practice during simulation, we treat this prefactor as a tunable parameter, calibrating the variance of the NN weights to ensure the field satisfies the canonical normalization of the free boson.
Emergence of the Virasoro Algebra.— The Virasoro generators Belavin:1984vu satisfy $L_{m},L_{n}]=(m-n)L_{m+n}+\frac{c}{12}m(m^{2}-1)\delta_{m+n,0}$. In the NN context, the generators $L_{n}(\Theta)$ are stochastic functionals of the random weights ($\Theta$). Expanding the network field into angular modes on the unit circle $z=e^{i\theta}$, the holomorphic current $J(z)=i\partial_{z}\phi$ decomposes into Laurent modes $\alpha_{m}(\Theta)$, which are specific Bessel-weighted sums of the network parameters (see SM for explicit derivation). The Virasoro generators then take the form of bilinears:
$L_{n}(\Theta)=\frac{1}{2}\sum_{m\in\mathbb{Z}}:\alpha_{n-m}(\Theta)\alpha_{m}(\Theta):\,$, (4)
where normal ordering in the NN context is defined by $AB:\equiv AB-\mathbb{E}[AB]$. The algebra is encoded in the statistics of the ensemble of randomly initialized network weights. Since the modes $\alpha_{m}$ are linear combinations of Gaussian weights, they satisfy $\mathbb{E}[\alpha_{m}\alpha_{n}]=mC_{\alpha\alpha}\delta_{m+n,0}$. We define the “neural bracket” via the connected correlation of the composite variables. Applying Isserlis’ theorem to the product $L_{m}L_{n}$:
$\mathbb{E}[L_{m}L_{n}]=\frac{1}{4}\sum_{j,k}\mathbb{E}[:\alpha_{m-j}\alpha_{j}::\alpha_{n-k}\alpha_{k}:]\,.$ (5)
This expectation decomposes into two contraction channels. First, contracting one pair of modes leaves a quadratic term proportional to $\alpha_{m+n-k}\alpha_{k}::$
$2\sum_{k}\mathbb{E}[\alpha_{m-k}\alpha_{n+k}]:\alpha_{k}\alpha_{n-k}:\longrightarrow(m-n)C_{\alpha\alpha}L_{m+n}\,.$ (6)
Crucially in the NN context, the amplitude of the generators depends on the variance of the network weights, $\sigma_{w}^{2}$, which sets the mode covariance scale $C_{\alpha\alpha}$. So the quadratic part of the algebra scales as $L_{n},L_{m}]\sim C_{\alpha\alpha}^{2}$, while the linear term scales as $C_{\alpha\alpha}$. Thus, the algebra closes with the correct normalization only if $C_{\alpha\alpha}=1$. This is the neural equivalent of enforcing Ward identities on the field. We calibrate the network hyperparameters to the $C_{\alpha\alpha}=1$ point.
Second, contracting all modes in pairs yields the vacuum variance:
$\frac{1}{2}\sum_{k}\mathbb{E}[\alpha_{m-k}\alpha_{k}]\mathbb{E}[\alpha_{k}\alpha_{m-k}]=\frac{C_{\alpha\alpha}^{2}}{2}\sum_{k}k(m-k)\,.$ (7)
This divergent sum, regularized by the spectral cutoff, yields the canonical anomaly term $\frac{c}{12}m(m^{2}-1)$ with $c=C_{\alpha\alpha}^{2}$. This finite-size scaling of the vacuum energy is the hallmark of conformal invariance on the cylinder Blote:1986qm . Thus, the full Virasoro algebra emerges from the neural modes. By tuning to the calibrated theory($C_{\alpha\alpha}=1$), we simulate a theory with $c=1$. We performed a numerical experiment to validate that the LK cos-net system describes the free scalar. The measured central charge was $c_{exp}=0.9787\pm 0.0421$; achieving agreement with the exact $c=1$ within $2.2\%$ NNFT_Repo .
Vertex Operators and Spectrum.— The “primary fields” of the free boson theory are Neural Vertex Operators (NVOs), defined not as operators on a Hilbert space, but as composite random variables on the network ensemble. We define an NVO of charge $a$ via the non-linear readout $\mathcal{V}_{a}(z;\Theta)\equiv:e^{ia\phi(z;\Theta)}:$. Normal ordering corresponds to normalization by the vacuum expectation $\mathbb{E}[e^{ia\phi}]$. Since the field $\phi$ is a Gaussian process in the infinite-width limit, the two-point function is evaluated exactly by averaging over the parameter priors using the identity for zero-mean variables $\mathbb{E}[e^{A}e^{B}]=e^{\frac{1}{2}\mathbb{E}[(A+B)^{2}]}$ :
$\mathbb{E}[\mathcal{V}_{a}(z)\mathcal{V}_{-a}(0)]$ $=\exp\left(a^{2}\mathbb{E}[\phi(z)\phi(0)]\right)$
$=|z|^{-2a^{2}}\,.$ (8)
This power-law decay identifies $\mathcal{V}_{a}$ as a conformal primary with scaling dimension, for the calibrated ensemble ($C_{\phi\phi}=1$), $\Delta_{a}=a^{2}$ (or $h=\bar{h}=a^{2}/2$). Our simulations reported in Fig. S1 confirm these exponents to $1\%$ for $a\leq 1$, demonstrating that the network captures the continuous spectrum of the free boson.
The Neural Majorana Fermion.— Following Frank:2025zuk , we define the network parameters for the fermionic theories as generators of a Grassmann algebra $\{\xi_{s}\}_{s=1}^{2N}$ satisfying $\{\xi_{s},\xi_{r}\}=0$. The ensemble is defined with respect to a symplectic structure $\Omega=\operatorname{diag}(\epsilon,\ldots,\epsilon)$ with symplectic unit $\epsilon=i\sigma_{2}$, so that $\mathbb{E}[\xi_{r}\xi_{s}]=(\Omega^{-1})_{rs}=-\Omega_{rs}$ and $\mathbb{E}[\xi_{2r-1}\xi_{2r}]=-\mathbb{E}[\xi_{2r}\xi_{2r-1}]=1$. The prior distribution is then
$P(\Theta)=e^{-\frac{1}{2}\xi^{T}\Omega\xi}$ (9)
where the Grassmann measure has preferred ordering $\int\mathrm{D}\xi:=\int\mathrm{d}\xi_{2N}\ldots\mathrm{d}\xi_{1}$ and $\int\mathrm{D}\xi P(\Theta)=\operatorname{Pf}(\Omega)=1$. All expectations $\langle\mathcal{O}\rangle=\mathbb{E}_{\xi}[\mathcal{O}]$ are defined by the Berezin integral against $\mathrm{D}\xi P(\Theta)$.
We define the Neural Majorana field (NMF)
$\psi(z)=N^{-1/2}\sum_{r=1}^{N}(\xi_{2r-1}u_{r}(x)+\xi_{2r}\mathcal{V}_{r}(z))$ (10)
weighted by holomorphic basis functions
$\begin{split}u_{n}(z)&=\sqrt{2k_{n}}e^{-i\theta_{k_{n}}}\cos(k_{n}\cdot z)~,\\ \mathcal{V}_{n}(x)&=i\sqrt{2k_{n}}e^{-i\theta_{k_{n}}}\sin(k_{n}\cdot z)~.\end{split}$ (11)
By exploiting the i.i.d. draws for the parameters and the $\xi_{r}$ algebra, the NMF propagator converges to the Cauchy-Kernel (CK):
$\mathbb{E}_{\xi}[\psi(z)\psi(w)]$ $=\frac{1}{N}\sum_{r=1}^{N}(u_{r}(z)\mathcal{V}_{r}(w)-\mathcal{V}_{r}(z)u_{r}(w))$
$\underset{N\to\infty}{=}\frac{1}{z-w}.$ (12)
The fermionic stress tensor is defined with respect to the neural normal ordering $T_{F}=-\frac{1}{2}:\psi(z)\partial\psi(z):$. Thus in the CK limit, by Wick’s theorem we find
$\mathbb{E}_{\xi}[T(z)T(w)]$ $=\frac{1}{4}\big(\mathbb{E}_{\xi}[\psi(z)\partial\psi(w)]\mathbb{E}_{\xi}[\partial\psi(z)\psi(w)]$
$\qquad\quad-\mathbb{E}_{\xi}[\psi(x)\psi(w)]\mathbb{E}_{\xi}[\partial\psi(z)\partial\psi(w)]$
$=\frac{1}{4(z-w)^{4}}.$ (13)
Hence, the NMF correctly realizes the $c=1/2$ free Majorana fermion. This identifies the NMF as the continuum limit of the free fermion sector of the 2d critical Ising model Schultz:1964fv . We leave the question of realizing the other primary fields of the critical Ising model and thus demonstrating that NNs can naturally represent minimal models with non-trivial statistics for future work.
Neural Dirac Fermion and Bosonization.— We construct a complex Neural Dirac Fermion $\Psi(z)=\frac{1}{\sqrt{2}}(\psi_{1}(z)+i\psi_{2}(z))$ from two independent NMFs. We define the $U(1)$ current $J_{F}=:\Psi^{\dagger}\Psi:=:i\psi_{1}\psi_{2}:$. Since the ensembles are independent, the current-current correlator factorizes:
$\mathbb{E}[J_{F}(z)J_{F}(w)]$ $=-\mathbb{E}[\psi_{1}(z)\psi_{1}(w)]\mathbb{E}[\psi_{2}(z)\psi_{2}(w)]$
$=\frac{1}{(z-w)^{2}}\,.$ (14)
This matches the bosonic current correlator $\langle i\partial\phi\,i\partial\phi\rangle\sim(z-w)^{-2}$ derived from the Log-Kernel. Thus, we see that two distinct architectures–LK boson and CK Dirac fermion–flow to the same universality class and generate identical current algebras, a strong signal that NN-FTs naturally realize bosonization Coleman:1974bu ; Mandelstam:1975hb and may offer a new framework for exploring other non-perturbative dualities such as mirror symmetry Hori:2000kt .
Emergence of the super-Virasoro Algebra.— With the formalism showing the emergence of Virsaoro symmetry from ensemble statistics in bosonic and fermionic theories, we are now in a position to derive the super-Virasoro algebra directly from the statistics of the neural modes. We realize this symmetry by constructing the holomorphic supercurrent $G(z)$ of a free $\mathcal{N}=(1,1)$ scalar multiplet; the construction of the anti-holomorphic current follows similarly. The field content of the scalar multiplet consists of the LK boson $\phi(z)$ and the NMF $\psi(z)$.
We decompose the ‘primary’ fields into modes $J(z)=\sum\alpha_{n}z^{-n-1}$ and $\psi(z)=\sum\beta_{r}z^{-r-1/2}$. Following from above, the statistics derived from the independent ensembles realize the mode algebra in the bosonic and fermionic sectors
$\begin{split}\mathbb{E}[\alpha_{n}\alpha_{m}]&=n\delta_{n+m,0}~,\\ \mathbb{E}[\beta_{r}\beta_{s}]&=\delta_{r+s,0}~.\end{split}$ (15)
We construct supercurrent as a composite $G=\psi J$ and decompose the stress tensor into a combination of bosonic and fermionic generators $L^{B}+L^{F}$. The mode expansions for these fields are given by Friedan:1985ge
$G_{r}$ $=\sum_{n\in\mathbb{Z}}\alpha_{n}\beta_{r-n}$ (16)
$L^{B}_{k}$ $=\frac{1}{2}\sum_{m\in\mathbb{Z}}:\alpha_{k-m}\alpha_{m}:$ (17)
$L^{F}_{k}$ $=\frac{1}{4}\sum_{q\in\mathbb{Z}+\frac{1}{2}}(2q-k):\beta_{k-q}\beta_{q}:$ (18)
Following the computations showing the emergence of the algebra of Virasoro generators, we evaluate the anti-commutator $\{G_{r},G_{s}\}$ via the symmetric expectation. In the product,
$G_{r}G_{s}=\sum_{n,m}\alpha_{n}\beta_{r-n}\alpha_{m}\beta_{s-m}=\sum_{n,m}(\alpha_{n}\alpha_{m})(\beta_{r-n}\beta_{s-m})\,$ (19)
there are three non-trivial contractions that need to be evaluated. First, the fermionic contraction $\mathbb{E}[\beta\beta]$ produces the bosonic generator: Using (15), the contraction forces m=r+snm=r+s-n, so
$\mathcal{C}_{\beta\beta}=\sum_{n}:\alpha_{n}\alpha_{r+s-n}:\mathbb{E}[\beta_{r-n}\beta_{-(r-n)}]=2L^{B}_{r+s}\,.$ (20)
Second, the bosonic contraction $\mathbb{E}[\alpha\alpha]$ produces the fermionic generator. Again using (15),
$\mathcal{C}_{\alpha\alpha}=\sum_{q}(q-s):\beta_{k-q}\beta_{q}:\,.$ (21)
Exploiting $\beta_{k-q}\beta_{q}:=-:\beta_{q}\beta_{k-q}:$ to symmetrize the sum yields
$\mathcal{C}_{\alpha\alpha}$ $=\frac{1}{2}\sum_{q}\left[(q-s)-(k-q-s)\right]:\beta_{k-q}\beta_{q}:$
$=2L^{F}_{r+s}\,.$ (22)
Finally, contracting all pairs forces $r+s=0$
$\mathcal{C}_{\alpha\alpha\beta\beta}$ $=\delta_{r+s,0}\sum_{n}\mathbb{E}[\alpha_{n}\alpha_{-n}]\mathbb{E}[\beta_{r-n}\beta_{-r+n}]$
$=\delta_{r+s,0}\sum_{n}^{r-1/2}n\,.$ (23)
The resulting sum yields the central extension $\frac{c}{3}(r^{2}-1/4)$ with $c=3/2$. Thus, we recover the full super-Virasoro algebra:
$\{G_{r},G_{s}\}=2L_{r+s}+\frac{c}{3}\left(r^{2}-\frac{1}{4}\right)\delta_{r+s,0}\,.$ (24)
Boundary NN-FTs.— We can extend our framework to the upper half plane $\mathbb{H}=\{z\in\mathbb{C}\mid\text{Im}(z)>0\}$ with boundary at $y=0$. Without using the embedding space formalism to introduce a boundary Capuozzo:2025ozt , we will use a more direct approach. Borrowing from the study of surface critical behavior Cardy:1984bb and quantum impurities Affleck:1995ge , we enforce conformal boundary conditions (BCs) via the method of images on the random features.
For the scalar field, we define the boundary field $\varphi_{\partial}$ by pairing each random feature $\phi_{j}(z)$ with a reflected image $\sigma\phi_{j}(\bar{z})$:
$\varphi_{\partial}(z)=\frac{1}{\sqrt{2N}}\sum_{j=1}^{N}\left[\phi_{j}(z)+\sigma\phi_{j}(\bar{z})\right]\,.$ (25)
Here, $\sigma=-1$ corresponds to Dirichlet BCs ($\varphi_{\partial}|=0$) and $\sigma=+1$ to Neumann BCs ($\partial_{y}\varphi_{\partial}|=0$). These BCs enforce $\partial_{x}\phi|=0$ or $\partial_{y}\phi|=0$, which implies that $\partial\phi|=\pm\bar{\partial}\bar{\phi}|$. Thus, the Cardy condition $T(z)|=\bar{T}(\bar{z})|$ holds, preserving the conformal subalgebra generated by $L_{n}+\bar{L}_{-n}$. Averaging over the ensemble yields the Boundary LK
$K(z,w)=-\ln|z-w|^{2}-\sigma\ln|z-\bar{w}|^{2}$ (26)
For the NMF $\psi$, the BCs couple the holomorphic field $\psi(z)$ to the anti-holomorphic field $\bar{\psi}(\bar{z})$ via a spin structure parameter $\eta=\pm 1$. We implement this by coupling the basis functions:
$\psi_{\partial}(z)=\frac{1}{\sqrt{2N}}\sum_{j=1}^{N}\left[\xi_{2j-1}u_{j}(z)+\eta\xi_{2j}\bar{u}_{j}(\bar{z})\right]\,.$ (27)
Averaging over the Grassmann ensemble yields the Boundary CK, where the reflection term is weighted by the spin structure:
$S_{B}(z,w)\equiv\mathbb{E}[\psi_{\partial}(z)\psi_{\partial}(w)]=\frac{1}{z-w}+\eta\frac{1}{z-\bar{w}}\,.$ (28)
We can now construct the full $\mathcal{N}=1$ Super-Kernel. We define the superfield $\Phi(Z)=\phi(z)+\theta\psi(z)$ in superspace coordinates $Z=(z,\theta)$. The super-propagator is $\mathcal{K}(Z_{1},Z_{2})=\langle\Phi(Z_{1})\Phi(Z_{2})\rangle=K_{B}+\theta_{1}\theta_{2}S_{B}$. Substituting the explicit kernels derived above, we find that the boundary terms combine into a super-invariant form only if the reflection parities match, $\eta=\sigma$:
$\mathcal{K}(Z_{1},Z_{2})=-\log|Z_{12}|^{2}-\sigma\log|Z_{1}-\bar{Z}_{2}|^{2}\,$, (29)
where $Z_{12}=z_{1}-z_{2}-\theta_{1}\theta_{2}$ is the supersymmetric interval. Expanding $\log(z-\theta_{1}\theta_{2})\simeq\log z-\tfrac{\theta_{1}\theta_{2}}{z}$ recovers the bosonic and fermionic propagators
$\mathcal{K}(Z_{1},\bar{Z}_{2})$ $=\left[-\ln|z_{12}|^{2}-\sigma\ln|z_{1}-\bar{z}_{2}|^{2}\right]$
$+\theta_{1}\theta_{2}\left[\frac{1}{z_{12}}+\sigma\frac{1}{z_{1}-\bar{z}_{2}}\right]\,$, (30)
where $z_{12}=z_{1}-z_{2}$. This explicitly demonstrates that preserving $\mathcal{N}=1$ supersymmetry requires imposing super-Neumann ($\sigma=\eta=+1$) or super-Dirichlet ($\sigma=\eta=-1$) BCs.
This architecture naturally realizes the Neveu-Schwarz (NS) sector. The Ramond (R) sector is accessible by introducing a $z^{1/2}$ twist
$\psi_{R}(z)$ $=z^{-1/2}\psi_{NS}(z)$ (31)
$S_{R}(z,w)$ $=\sqrt{z^{-1}w^{-1}}S_{B}(z,w)$ (32)
introducing the branch cuts necessary for the Ramond ground state.
Discussion.— We have established a framework for constructing local 2d CFTs from NNs, identifying the LK architecture as a scalar Gaussian fixed point with Virasoro symmetry. By utilizing non-linear readouts and the method of images, we further demonstrated that NN-FTs can simulate theories with fermionic statistics and boundaries.
While the results of this work have rigorously defined a free fixed point ($c=1$) for an NN-FT realizing a 2d bosonic theory in the infinite width limit, the framework naturally includes interactions furnished by finite width effects. In SM, we provide experimental evidence obtained by simulating $\langle\phi\phi\phi\phi\rangle$ using up to $N=512$ Fourier features at the input layer where we observe that the excess kurtosis of the field distribution is non-zero Halverson:2020trp and scales as $1/N$. This establishes the $N\to\infty$ limit as a definition of the ‘perturbative vacuum’ for these NN-FTs. We will leave the theoretical analysis of these finite width corrections Yaida:2020 ; Roberts:2021principles and the engineering of interacting 2d conformal theories for future work.
Acknowledgements.— We thank Jim Halverson, Benjamin Suzzoni, and Pietro Capuozzo for invaluable comments on the draft. The work of B.R. is supported by an NWO vidi grant (number 016.Vidi.189.182).
Supplemental Material for Virasoro Symmetry in Neural Network Field Theories
Appendix A S1. Uniqueness of the Log-Kernel
In this section, we provide a proof that the scale-invariant spectral density $p(k)\propto|k|^{-2}$ is the unique choice for a rotation-invariant network prior that yields a well-defined Virasoro algebra in the infinite width limit of a 2d NN-FT describing a free boson.
Consider the infinite-width limit of the NN $\phi(x)$, which defines a Gaussian Process. For the theory to admit a local conformal structure, the stress tensor $T(z)\sim:(\partial\phi)^{2}:$ must have a two-point function decaying as $z^{-4}$. This requires the gradient field $J(z)=i\partial_{z}\phi(x)$ to have a covariance scaling as:
$\mathbb{E}[J(z)J(w)]\propto\frac{1}{(z-w)^{2}}\,.$ (S1)
We assume a generic rotationally invariant prior $p(k)\propto|k|^{-\alpha}$. The covariance of the gradient $\partial\phi$ is given by the Fourier transform of the power spectrum weighted by the momentum squared (from the derivatives):
$G_{\partial\phi}(r)=\int d^{2}k\,|k|^{2}p(k)e^{ik\cdot x}\propto\int_{0}^{\infty}dk\,k\cdot k^{2}\cdot k^{-\alpha}\int_{0}^{2\pi}d\theta\,e^{ikr\cos\theta}\,.$ (S2)
The angular integral yields the Bessel function $2\pi J_{0}(kr)$. The radial integral becomes:
$G_{\partial\phi}(r)\propto\int dk\,k^{3-\alpha}J_{0}(kr)\,.$ (S3)
By dimensional analysis, this integral scales as $r^{-(4-\alpha)}$. To match the conformal requirement $G(r)\sim r^{-2}$, we must have $4-\alpha=2$, which uniquely fixes the spectral exponent to $\alpha=2$}. Any other choice of $\alpha$ would result in a stress tensor with anomalous scaling dimensions, violating the Virasoro algebra.
Appendix B S2. Derivation of Neural Modes
We explicitly derive the relation between the random Fourier features and the Laurent modes of the conformal field.
B.1 Bosonic Modes
The network output is defined as $\phi(x)=\frac{1}{\sqrt{N}}\sum_{j}\cos(k_{j}\cdot x+\gamma_{j})$. We expand the cosine using the Jacobi-Anger identity:
$e^{ik\cdot x}=e^{i|k||x|\cos(\theta_{k}-\theta_{x})}=\sum_{n\in\mathbb{Z}}i^{n}J_{n}(|k|r)e^{in(\theta_{x}-\theta_{k})}\,.$ (S4)
Substituting this into the network definition and isolating the term transforming as $e^{-im\theta_{x}}$ (the $m$-th angular momentum mode), we find the expansion for the holomorphic current modes $\alpha_{m}$. Specifically, requiring $J(z)=\sum\alpha_{m}z^{-m-1}$ implies that $\alpha_{m}$ extracts the radial dependence $r^{-m-1}$. The final expression for the stochastic mode functional is:
$\alpha_{m}(\Theta)=-\frac{1}{4\sqrt{N}}\sum_{j=1}^{N}|k_{j}|r^{m+1}\Big(i^{-m-1}J_{-m-1}(|k_{j}|r)e^{i((m+1)\theta_{k_{j}}+\gamma_{j})}+\text{h.c.}\Big)\,.$ (S5)
While this expression depends explicitly on $r$, the ensemble expectation $\mathbb{E}[\alpha_{m}\alpha_{n}]$ involves an integral $\int dk\,k^{-1}J_{\nu}(kr)J_{\mu}(kr)$ which becomes independent of $r$ for the Log-Kernel distribution, confirming the conformal invariance of the modes.
B.2 Fermionic Modes and Spin Structures
B.2.1 Neveu-Schwarz (NS) Sector
The standard neural ansatz $\psi(z)$ uses basis functions $e^{ikz}$ which are single-valued on the complex plane. The Jacobi-Anger expansion yields modes with integer angular momentum $n\in\mathbb{Z}$. Comparing this to the Laurent expansion for a dimension $1/2$ field:
$\psi(z)=\sum_{r}\beta_{r}z^{-r-1/2}=\sum_{r}\beta_{r}r^{-r-1/2}e^{-i(r+1/2)\theta}\,.$ (S6)
Matching the angular dependence n=r1/2n=-r-1/2 implies r=n1/2r=-n-1/2. Since $n$ is an integer, the mode indices $r$ must be half-integers ($r\in\mathbb{Z}+1/2$). This identifies the bulk architecture with standard single-valued neurons as the Neveu-Schwarz (NS) sector. The explicit mode form is:
$\beta_{r}^{\text{NS}}=\frac{1}{\sqrt{N}}\sum_{j}\xi_{j}\sqrt{k_{j}}(-ir)^{r+1/2}J_{-r-1/2}(k_{j}r)e^{i((r+1/2)\theta_{k_{j}}+\gamma_{j})}\,.$ (S7)
B.2.2 Ramond (R) Sector and the Zero Mode
The Ramond sector requires periodic boundary conditions on the cylinder $\psi(w+2\pi i)=+\psi(w)$, which maps to a branch cut on the plane $\psi_{R}(z)=z^{-1/2}\psi_{\text{NS}}(z)$. Expanding the twisted field yields integer modes $\varsigma_{n}\in\mathbb{Z}$:
$\psi_{R}(z)=\sum_{n\in\mathbb{Z}}\varsigma_{n}z^{-n-1/2}\,.$ (S8)
The twist field maps the half-integer index $r$ of the NS sector to the integer index $n=r+1/2$. For non-zero modes ($n\neq 0$), we derive the explicit neural representation by substituting r=n1/2r=n-1/2 into the NS result (Eq. S7):
$\varsigma_{n}(\Theta)=\frac{1}{\sqrt{N}}\sum_{j=1}^{N}\xi_{j}\sqrt{k_{j}}(-ir)^{n}J_{-n}(k_{j}r)e^{i(n\theta_{k_{j}}+\gamma_{j})}\quad(n\neq 0)\,.$ (S9)
These modes satisfy the standard fermionic anti-commutation relations $\mathbb{E}[\varsigma_{n}\varsigma_{m}]=\delta_{n+m,0}$.
The zero mode $\varsigma_{0}$, however, requires special treatment. In a generic random feature network, $\varsigma_{0}$ would be a Grassmann number with $\varsigma_{0}^{2}=0$. To realize the Ramond vacuum degeneracy, $\varsigma_{0}$ must satisfy the Clifford algebra $\{\varsigma_{0},\varsigma_{0}\}=1\implies\varsigma_{0}^{2}=1/2$. We achieve this by promoting the network readout for the zero-momentum component to a matrix-valued operation acting on an internal 2d spin space (representing the vacuum states $|\sigma_{\pm}\rangle$). We assign the zero-mode feature a Clifford weight:
$\varsigma_{0}(\Theta)=\frac{1}{\sqrt{2}}\gamma^{5}\otimes\mathds{1}_{\text{Grassmann}}\,.$ (S10)
Under this construction, the zero mode squares to the identity on the spin space, $\varsigma_{0}^{2}=1/2\cdot\mathbb{I}$, recovering the exact quantum algebra necessary for the superconformal ground state.
Appendix C S3. Numerical Validations
C.1 Central Charge and Normal Ordering
We determined the central charge $c$ by extracting the coefficient of the singular term in the stress tensor OPE, $\langle T(z)T(0)\rangle=\frac{c}{2z^{4}}$. Directly simulating this correlator is numerically unstable due to the large vacuum energy of the gradient field, which scales as $\Lambda_{UV}^{2}$. To mitigate this, we employed a two-pass variance reduction algorithm to enforce normal ordering at the ensemble level.
In the first pass, we simulate a pilot ensemble of $M_{cal}=10^{4}$ networks to determine the vacuum expectation value of the squared current, $\sigma_{J}^{2}\equiv\mathbb{E}[J(z)^{2}]$. Due to the stationarity of the Log-Kernel, this value is position-independent. This defines the numerical vacuum energy.
In the second pass, we then simulated the production ensemble of $M=5\times 10^{5}$ networks with width $N=10^{4}$. For each realization $\Theta_{k}$, we constructed the normal-ordered stress tensor by subtracting the pre-computed vacuum energy:
$T(z;\Theta_{k})=-\frac{1}{2}\left(:J(z)^{2}:\right)\equiv-\frac{1}{2}\left(J(z;\Theta_{k})^{2}-\sigma_{J}^{2}\right)\,.$ (S11)
The correlation function was then estimated via the ensemble average:
$\hat{G}_{TT}(r)=\frac{1}{M}\sum_{k=1}^{M}T(r;\Theta_{k})T(0;\Theta_{k})\,.$ (S12)
This subtraction technique reduces the variance of the estimator by orders of magnitude compared to computing $\langle J^{2}(r)J^{2}(0)\rangle$ directly.
We computed $\hat{G}_{TT}(r)$ for 50 points $r\in[0.1,5.0]$. The central charge was extracted by fitting the data to the theoretical curve $f(r)=\frac{A}{r^{4}}$ in the window $r\in[0.5,3.0]$. The fitted amplitude yielded $A=c_{exp}/2$. For the calibrated ensemble (where $C_{\phi\phi}$ was tuned to unity), we obtained $c_{exp}=0.9787\pm 0.0421$, which agrees with the exact value $c=1$ within $2.2\%$.
C.2 Vertex Operator Spectrum Experiment
To verify the continuous spectrum of scaling dimensions, we constructed Neural Vertex Operators (NVOs) $\mathcal{V}_{\alpha}(z)=:e^{i\alpha\phi(z)}:$ for various charges $\alpha$. The scaling dimension $\Delta_{\alpha}$ was extracted from the power-law decay of the two-point function $G_{\alpha}(r)=\langle\mathcal{V}_{\alpha}(0)\mathcal{V}_{-\alpha}(r)\rangle$.
We simulated an ensemble of $M=5\times 10^{5}$ networks with width $N=5000$. For each realization $\Theta_{k}$, we computed the field at the origin $\phi(0)$ and at 20 radial points $r_{j}$ logarithmically spaced between $r_{min}=0.1$ and $r_{max}=10.0$ (in units of $\Lambda_{UV}^{-1}$). The correlation function was estimated via the Monte Carlo average:
$\hat{G}_{\alpha}(r_{j})=\frac{1}{M}\sum_{k=1}^{M}\exp\left[i\alpha\left(\phi(r_{j};\Theta_{k})-\phi(0;\Theta_{k})\right)\right]\,.$ (S13)
Note that we use the difference $\phi(r)-\phi(0)$ to automatically enforce charge neutrality and cancel the zero-mode divergence, removing the need for explicit normal ordering subtractions in the simulation.
Figure S1: Log-log plot of the Neural Vertex Operator two-point function. The slopes correspond to the conformal scaling dimensions $2\Delta$. The dashed lines indicate the theoretical prediction $\Delta=\alpha^{2}$ for the free boson, showing excellent agreement with the neural simulation.
Theory predicts $G_{\alpha}(r)\sim r^{-2\Delta_{\alpha}}$. We extracted $\Delta_{\alpha}$ via linear regression of $\ln|\hat{G}_{\alpha}(r)|$ against $\ln r$. To avoid cutoff artifacts, the fit was restricted to the physical window $r\in[0.5,5.0]$. We performed this measurement for charges $\alpha\in\{0.5,1.0,1.5,2.0\}$. The extracted exponents were compared to the free boson prediction $\Delta_{th}=\alpha^{2}$ (assuming $C_{\phi\phi}=1$). The agreement was robust: for $\alpha=1$, we measured $\Delta=1.012\pm 0.008$. For higher charges, the variance of the estimator grows exponentially; nonetheless, we recovered $\Delta=4$ for the $\alpha=2$ operator within $3.5\%$ accuracy.
C.3 Gaussianity and Finite-Width Interactions
To verify that the Log-Kernel architecture converges to the free boson fixed point, and to quantify the strength of finite-width interactions, we analyzed the connected four-point function $G_{4c}$. In a scalar field theory, this correlator determines the effective coupling constant $\lambda$ of the interaction term $\frac{\lambda}{4!}\phi^{4}$.
We measured the excess kurtosis of the field at a single point. For an ensemble of size $M$, the unbiased estimator for the connected component is:
$G_{4c}\approx\hat{\mu}_{4}-3(\hat{\mu}_{2})^{2}\,$, (S14)
where $\hat{\mu}_{n}=\frac{1}{M}\sum_{k=1}^{M}\phi(0;\Theta_{k})^{n}$ are the raw sample moments. For a Gaussian distribution, this quantity vanishes exactly. In our effective theory, we predict a scaling $|G_{4c}|\sim C/N$, where $C$ is an architecture-dependent constant and $N$ is the network width.
To resolve the signal against Monte Carlo noise, we required an extremely large ensemble. We simulated $M=10^{8}$ independent networks for widths $N$ logarithmically spaced from $N=2$ to $N=512$. The computations were performed in batches to manage memory constraints. We plotted the magnitude $|G_{4c}|$ versus $N$ on a log-log scale.
The results (Fig. S2) reveal two distinct regimes:
  1. 1.
    Perturbative Regime ($N<256$): The data follows a precise power law. A linear regression gives a slope of $1.02\pm 0.01$, confirming the theoretical prediction of $\mathcal{O}(1/N)$ scaling for the interactions.
  2. 2.
    Noise Regime ($N\geq 512$): The signal flattens as it hits the Monte Carlo noise floor $\sim 1/\sqrt{M}\approx 10^{-4}$.
This confirms that the infinite-width limit is a trivial Gaussian fixed point, and that non-trivial interactions are systematically generated by finite-$N$ corrections. The high precision of the slope in the perturbative regime provides strong numerical evidence for the validity of the $1/N$ expansion in this architecture.
Figure S2: Scaling of the connected four-point function $G_{4c}$ (interactions) with network width $N$. The blue line represents simulation data ($M=10^{8}$), which closely tracks the theoretical $1/N$ scaling (red dashed line). The signal is clearly resolvable in the perturbative regime before hitting the statistical noise floor at $N\approx 512$.
Appendix D S4. Simulation Pseudocode
To facilitate reproducibility, we provide the explicit algorithms used to generate the numerical results. The source code is available at NNFT_Repo .
D.1 Algorithm 1: Central Charge (Two-Pass Method)
This algorithm implements the vacuum subtraction strategy to compute $c$. The function GetCurrent(z) computes $J(z)=i\partial\phi(z)$ for a random initialization.
# Parameters
M_cal = 10^4       # Calibration size
M_sim = 5*10^5     # Simulation size
N = 10^4           # Network width
r_vals = [0.1 ... 5.0]
# PASS 1: Calibrate Vacuum Energy
vac_energy = 0
for k in 1 to M_cal:
    Theta = InitializeWeights(N)
    J_0 = GetCurrent(0, Theta)  # Sample current at origin
    vac_energy += (J_0 * J_0) / M_cal
# PASS 2: Measure Correlator
G_TT = zeros(len(r_vals))
for k in 1 to M_sim:
    Theta = InitializeWeights(N)
    # Compute Normal Ordered T at 0 and r
    J_0 = GetCurrent(0, Theta)
    T_0 = -0.5 * (J_0^2 - vac_energy)
    for r in r_vals:
        J_r = GetCurrent(r, Theta)
        T_r = -0.5 * (J_r^2 - vac_energy)
        # Accumulate Correlator
        G_TT[r] += (T_0 * T_r) / M_sim
# Extract c via regression on G_TT ~ c / (2*r^4)
c_exp = FitPowerLaw(r_vals, G_TT) * 2
D.2 Algorithm 2: Vertex Operator Scaling
This algorithm computes the scaling dimension $\Delta_{\alpha}$ using the difference method to enforce charge neutrality.
# Parameters
alpha = 1.0        # Charge
M = 5*10^5         # Ensemble size
r_vals = LogSpace(0.1, 10.0, 20)
G_V = zeros(len(r_vals))
for k in 1 to M:
    Theta = InitializeWeights(N)
    phi_0 = GetField(0, Theta)
    for r in r_vals:
        phi_r = GetField(r, Theta)
        # Compute Vertex Operator product :e^ia phi(0): :e^-ia phi(r):
        # equivalent to e^(i*alpha*(phi(0) - phi(r)))
        phase = alpha * (phi_0 - phi_r)
        term = ComplexExp(1j * phase)
        G_V[r] += term / M
# Extract Delta via regression on log(G_V) ~ -2*Delta * log(r)
Delta = -0.5 * Slope(log(r_vals), log(G_V))
D.3 Algorithm 3: Finite-Width Interactions ($G_{4c}$)
This algorithm measures the interaction strength (connected kurtosis) as a function of network width.
# Parameters
Widths = [2, 4, 8, ... 512]
M = 10^8           # High precision ensemble
Results = []
for N in Widths:
    mu_2 = 0       # Second Moment
    mu_4 = 0       # Fourth Moment
    for k in 1 to M:
        Theta = InitializeWeights(N)
        phi_val = GetField(0, Theta)
        mu_2 += (phi_val^2) / M
        mu_4 += (phi_val^4) / M
    # Connected 4-point function (Wick subtraction)
    # G_4c = <phi^4> - 3<phi^2>^2
    G_4c = mu_4 - 3 * (mu_2^2)
    Results.append( (N, abs(G_4c)) )
# Verify 1/N scaling
Exponent = Slope(log(Widths), log(Results))